JNTU

JNTU B.Tech II PROBABILITY AND STATISTICS Previous Question Paper (December - 2018 - R13)

B.Tech | II | II-II | R13 | December - 2018

Answer PaperPaper Code: 114DKMaximum Marks: 75Time: 3 Hours

Note: This question paper contains two parts A and B. Part A is compulsory which carries 25 marks. Answer all questions in Part A. Part B consists of 5 Units. Answer any one full question from each unit. Each question carries 10 marks and may have a, b, c as sub questions.

Part A

(25 Marks)

125 marks
a

The probability density function is f(x) = K(3x^2 - 1) for -1 ≤ x ≤ 2, and 0 elsewhere. Find K.

2 marks

Answer

To find K, we use the condition that the total area under the probability density function must be 1:

∫-12 K(3x2 - 1) dx = 1

Integrating the function:

K [x3 - x]-12 = 1
K [(8 - 2) - (-1 + 1)] = 1
K [6 - 0] = 1

Thus, K = 1/6.

b

Six cards are drawn from a pack of 52 cards. Getting a red card is a success. Find the probability of getting the success at least once.

3 marks

Answer

The probability of drawing a red card in a single draw is p = 26/52 = 0.5. The probability of not drawing a red card is q = 0.5. For n = 6 trials, the probability of getting at least one success is 1 minus the probability of zero successes:

P(X ≥ 1) = 1 - P(X = 0) = 1 - (0.5)6
P(X ≥ 1) = 1 - 0.015625 = 0.984375
c

The joint probability density function is f(x, y) = (6-x-y)/8 for 0 < x < 2, 2 < y < 4, and 0 elsewhere. Find the Marginal probability density function for X.

2 marks

Answer

Sign in to view the complete answer paper.

d

The equations of two Regression lines are y=11.64-0.5x, x=19.13-0.87y, find the Coefficient of Correlation.

3 marks

Answer

Sign in to view the complete answer paper.

e

Define tail I and tail II tests.

2 marks

Answer

Sign in to view the complete answer paper.

f

A Random sample of 1500 items is taken from a population whose standard deviation is 45. Find the maximum error.

3 marks

Answer

Sign in to view the complete answer paper.

g

Define Waiting time and Mean arrival rate.

2 marks

Answer

Sign in to view the complete answer paper.

h

What are the assumptions for the postulates for the poisson process?

3 marks

Answer

Sign in to view the complete answer paper.

i

Define an irreducible Marcov chain.

2 marks

Answer

Sign in to view the complete answer paper.

j

What are the properties of a transition matrix?

3 marks

Answer

Sign in to view the complete answer paper.

Part B

(50 Marks)

210 marks
a

A die is thrown 8 times. If getting 2 or 4 is a success, find the probability of getting: i) 4 successes ii) P(x ≤ 3) (x being the number of failures)

5 marks

Answer

Sign in to view the complete answer paper.

b

Suppose that the weights of 800 male students are normally distributed with mean μ = 140 pounds with a standard deviation of 10 pounds. Find the number of students whose weight are: i) Between 138 and 148 ii) More than 152.

5 marks

Answer

Sign in to view the complete answer paper.

OR
310 marks
a

If the variance of a Poisson variate is 2, find the probability that: i) x =0 ii) 1 < x < 4.

5 marks

Answer

Sign in to view the complete answer paper.

b

If the masses of 300 students are normally distributed with mean 68 kgs and standard deviation 3 kgs. How many students have masses i) Greater than 72 kgs ii) Between 65 and 71 kgs.

5 marks

Answer

Sign in to view the complete answer paper.

4

The joint probability density function of the continuous random variables X and Y is given by f(x,y). Find f(x) and f(y).

10 marks

Answer

Sign in to view the complete answer paper.

OR
5

Calculate the coefficient of rank correlation for the provided X and Y data table.

10 marks

Answer

Sign in to view the complete answer paper.

6

In a city A 20% of a random sample of 900 school boys had a certain slight physical defect. In another city B 18.5% of a random sample of 1600 school boys had the same defect. Is the difference between the proportions significant at 0.05 level of significance.

10 marks

Answer

Sign in to view the complete answer paper.

OR
7

From the following data find whether there is any significant liking in the habit of taking soft drinks among the categories of employees (Table provided).

10 marks

Answer

Sign in to view the complete answer paper.

8

The milk plant at a city distributes its products by trucks. The trucks arrive at the interval of 20 minutes. The service time is 4 minutes. a) What is the probability that there are more than or equal to 4 trucks in the queue? b) What is the waiting time of a truck in the queue? c) What is the variance of queue length?

10 marks

Answer

Sign in to view the complete answer paper.

OR
9

Problems arrive at a computer centre in Poisson fashion with a mean arrival rate of 25 per hour. The average computing job requires 2 minutes of terminal time. Calculate the following: a) Average number of problems waiting for the queue. b) The probability that queue length is greater than or equal to five. c) The idle time.

10 marks

Answer

Sign in to view the complete answer paper.

10

There are three stores of groceries. A study found initial distribution and monthly transition probabilities. a) What proportion of customers retain by Feb 1st and March 1st b) Find the long run proportions.

10 marks

Answer

Sign in to view the complete answer paper.

OR
11

In a presidential election, initial voter registration is provided along with monthly transition probabilities. a) Write transition probability matrix b) Find the % of the voters in each category at the end July.

10 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the question paper of PROBABILITY AND STATISTICS from JNTU for B.Tech CE II II-II, December - 2018, R13.

For the solved question paper with answers, switch to the Answer Paper tab.

Other papers in PROBABILITY AND STATISTICS

Loading Question Papers

Fetching data from database...